How would I solve y = 15 - 3x and 7y + 3x = 15 using the substitution method? I tried plugging the y = 15 - 3x for the y but I came up with a decimal answer and I don't think it's supposed to be a decimal?
Answer by mananth(16946)

3 x + 1 y = 15 .............1
7 y = 15 -3 x ---------2
/ 7 =
y = 2.14 -0.43 x
Substitute y in (1)
3 x + 1 ( 2.14 -0.43 x ) = 15
3 x + 2.14 + -0.43 x = 15
3 x -0.43 x = 15 -2.14
2.57 x = 12.86
/ 2.57
x= 5
Plug the value of x in (1)
3 x + 1 y = 15
3 * 5 + 1 y = 15
15 + 1 y = 15
1 y = 0
/ 1
y= 0
m.ananth@hotmail.ca